# A Study of Optimizations for Fine-tuning Large Language Models

### v1

@@ Abstract
Fine-tuning large language models is a popular choice among users trying to adapt them for specific applications. However, fine-tuning these models is a demanding task because the user has to examine several factors, such as resource budget, runtime, model size and context length among others. A specific challenge is that fine-tuning is memory intensive, imposing constraints on the required hardware memory and context length of training data that can be handled. In this work, we share a detailed study on a variety of fine-tuning optimizations across different fine-tuning scenarios. In particular, we assess Gradient Checkpointing, Low Rank Adaptation, DeepSpeed’s ZeRO Redundancy Optimizer and Flash Attention. With a focus on memory and runtime, we examine the impact of different optimization combinations on GPU memory usage and execution runtime during fine-tuning phase. We provide recommendation on best default optimization for balancing memory and runtime across diverse model sizes. We share effective strategies for fine-tuning very large models with tens or hundreds of billions of parameters and enabling large context lengths during fine-tuning. Furthermore, we propose the appropriate optimization mixtures for fine-tuning under GPU resource limitations.
@@ 1 Introduction
The remainder of this work is organized as follows. We review the existing fine-tuning optimization literature in section 2. Section 3 presents an overview of commonly used fine-tuning optimizations: Gradient Checkpointing, Low Rank Adaptation, DeepSpeed ZeRO Redundancy Optimizer and Flash Attention. We then describe in Section 4 the theory behind GPU memory usage and compare theoretical values with our empirical results. In Section 5, we discuss our experimental results and analyze the impact of using fine-tuning optimizations in conjunction with each other. Finally, we share the key conclusions and directions for future work in Section 6.
@@ 2 Related Work
Fine-tuning large language models is a computational and memory intensive process. Several optimizations proposed in existing literature either reduce the memory consumption, or the runtime, or accomplish both during fine-tuning. Tensor and pipeline parallelism are types of model parallelism techniques which enable large models to fit in memory for training. In tensor parallelism (Shoeybi et al., 2020), the model is split tensor-wise among GPUs. Conversely, pipeline parallelism (Huang et al., 2019) distributes the model layer-wise across GPUs. DeepSpeed ZeRO (Rajbhandari et al., 2020) integrates data and model parallelism, by partitioning model states across data-parallel processes, which are reconstructed when required during training. Fully Sharded Data Parallel (FSDP) (FairScale authors, 2021) is similar to DeepSpeed ZeRO and supports model parameter sharding to train large models.
Contrary to parallelism strategies, Parameter-Efficient Fine-Tuning (PEFT) methods (Ding et al., 2023; Han et al., 2024) modify a small proportion of model parameters, while keeping the others frozen during training. This reduction in total trainable parameters subsequently lowers the memory footprint while still achieving comparable performance to full fine-tuning. Low Rank Adaptation (Hu et al., 2021) is a widely used PEFT technique, where low rank tensor formulation of model parameters are updated in the training process. Prefix tuning (Li and Liang, 2021) and Prompt tuning (Lester et al., 2021) are examples of PEFT method where trainable embeddings are injected at the model or input level, while the entire model is frozen. PEFT methods can easily integrate with parallelism techniques such as DeepSpeed ZeRO, enhancing the gains in terms of memory and runtime reduction during fine-tuning.
Certain optimizations are designed to address specific memory bottlenecks during fine-tuning. Gradient computation during backward pass can consume significant GPU memory for large models because of the need to maintain large number of activations in memory. Gradient Checkpointing (Sohoni et al., 2019) saves only few activations from forward pass, while recomputing most of them during backward pass, to reduce the memory usage at the cost of increased runtime. Attention calculations can form memory bottlenecks when context length is very large. Flash Attention (Dao et al., 2022; Dao, 2023) methods focus on optimizing the memory consumed during attention calculation operations. Both Gradient Checkpointing and Flash Attention can be combined with PEFT and DeepSpeed ZeRO.
@@ 3.1 Gradient Checkpointing
3.2 Low Rank Adaptation
Low Rank Adaptation (LoRA) reduces the number of trainable parameters which in turn lowers GPU memory requirements during fine-tuning. It achieves this by freezing the pre-trained model weights and subsequently injecting trainable rank decomposition matrices into each of the chosen layers of the Transformer architecture (Hu et al., 2021). The number of trainable parameters are reduced by orders of magnitude, bringing down the cost of fine-tuning, while also preserving the quality of the results. As an example, with LoRA fine-tuning using rank set to 64, the number of trainable parameters for a 70 billion parameter model is reduced to about 131 million parameters ($\sim$0.19of the original model size).
3.3 DeepSpeed
DeepSpeed’s ZeRO Redundancy Optimizer (ZeRO) (Rajbhandari et al., 2020) is a memory optimization technique that provides the benefits of model and data parallelism, while alleviating the limitations of both. ZeRO powered Data Parallelism (Zero-DP) partitions the model states – parameters, gradients, and optimizer states – across the data-parallel processes and uses a dynamic communication schedule to share the necessary model states across processes. ZeRO-DP provides three optimization stages, which successively provide larger and larger memory reduction while incurring some runtime overhead. ZeRO-DP based fine-tuning can achieve memory reduction from 4 to 8 times for Stages 1 and 2 respectively, and all the way up to linear memory reduction for Stage 3 (given a large number of GPUs). This does come at the expense of runtime, especially for ZeRO-DP Stage 3. ZeRO optimization is further enhanced by the inclusion of ZeRO-Offload and ZeRO-Infinity (Ren et al., 2021; Rajbhandari et al., 2021), which enable offloading the optimizer states and model parameters to CPU respectively. While ZeRO-Offload is available across all three ZeRO-DP stages, ZeRO-infinity is available only for stage 3. Without using DeepSpeed ZeRO kind of optimizations, many large models (several billions of parameters) are practically impossible to train using just a handful of V100 or A100 GPUs.
@@ 3.4 Flash Attention
Flash Attention helps achieve attention calculations in linear instead of quadratic complexity with respect to the sequence/context length (Dao et al., 2022). It leverages tiling and recomputation techniques to significantly speed up attention computation. It makes judicious use of the Static Random Access Memory (SRAM), which is the most expensive but smallest sized memory unit within the GPU, minimizing reads and writes between SRAM and HBM. In Flash Attention version 2, there are further optimizations to reduce the slower non-matrix multiplication operations and to parallelize forward and backward passes along the sequence length dimension (in addition to the batch and number of heads dimensions) (Dao, 2023).
@@ 4 Theoretical Analysis of GPU Memory Requirements
Table 1: Comparison between theoretical and empirical GPU memory allocated (in GB) across the three stages of ZeRO-DP. Fine-tuned model is LLaMA-2-7B using 8xA100 GPUs (each with 80 GB of HBM). The theoretical estimates can be used to approximate the empirical memory requirements before running the fine-tuning jobs.
Table 1 shows our results of comparing the theoretically expected and empirical GPU memory allocated for fine-tuning LLaMA-2-7B across ZeRO-DP stages. We use the notation ZeRO-1, ZeRO-2 and ZeRO-3 to denote ZeRO-DP stages 1, 2 and 3 respectively. The theoretical memory is computed by summing the memory from model states and activations. We use the following model default parameters: number of transformer layers=32, hidden dimension size $h$=4096 and attention heads $a$=32 to theoretically compute the total activation memory of 1.48 GB. The empirical GPU memory refers to the peak GPU memory allocated during fine-tuning. We show empirical memory with and without CPU offloading of optimizer states. The experimental setup details are described in subsection 5.1.
@@ 5 Experiments
We begin this section by analyzing the impact of fine-tuning optimizations on memory and runtime. The results of our analysis are used to recommend a set of optimizations that can act as a balanced default to optimize for memory and runtime during fine-tuning. Subsequently, we delve into fine-tuning large models (with tens of billions of parameters) and explore the optimizations which make such fine-tuning possible. Following this, we examine the role of Flash Attention 2 during fine-tuning of LLMs on long context data. Lastly, we investigate the specific case of fine-tuning with resource-constrained GPUs, specifically V100s. Our goal is to identify the fine-tuning optimizations that can enable efficient fine-tuning across various model sizes and context lengths, even with limited resources.
@@ 5.1 Setup
We fine-tune models from LLaMA-2 (7B, 13B, 70B) and Falcon (180B) families on Causal Language Modeling task. Following are the details of our experimental setup:
@@ 5.2 Best Default Optimizations to Balance Memory and Runtime
Figure 2: GPU memory usage and fine-tuning runtime for different optimization configurations across ZeRO-1, ZeRO-2 and ZeRO-3 for LLaMA-2 7B. Using LoRA with ZeRO-2 provides the best balance between memory usage and runtime.
We have already established that for a relatively small 7B parameter model, it is infeasible to fine-tune without any optimizations, even when using A100s. Therefore, we use DeepSpeed ZeRO which provides both model and data parallelism. We examine the set of fine-tuning optimizations that should be enabled by default in a fine-tuning framework. Our criteria for selecting such optimizations is based on their ability to strike an optimal balance between memory usage and runtime in conjunction with ZeRO-DP.
We fine-tune LLaMA-2-7B on one node (8xA100 GPUs) with five distinct configurations: (a) no optimization enabled (b) with gradient checkpointing (c) with LoRA (d) with flash attention 2 and (e) without CPU offloading. Each configuration is enabled with all three ZeRO-DP stages. We chose A100 GPUs for these experiments due to their substantial memory capacity of 80 GB. This ample memory allows us the flexibility to experiment with various optimizations by toggling them on and off, while avoiding out-of-memory failures.
@@ 5.3 Fine-tuning Large Models
Fine-tuning large models such as LLaMA-2 70B or Falcon 180B can often run into out-of-memory (OOM) errors if the right set of optimizations are not enabled. Enabling ZeRO-3 is a viable alternative in such scenarios though it comes at the expense of increased runtime. Let’s examine the model state memory term in ZeRO-3, fine-tuned over $M$ nodes with $N$ GPUs per node:
@@
LLaMa-2 Model GPU Memory Usage (GB) Time (seconds)
@@
Table 2: GPU memory usage and fine-tuning time for LLaMA-2 Models with ZeRO-3 + LoRA on 8xV100 GPUs.
We were able to fine-tune LLaMA-2 70B using a combination of ZeRO-3 + LoRA on a single node with 8xV100 GPUs (32 GB HBM). Table 2 shows the actual GPU memory consumed to be around 15.54 GB for the 70B model in our experiments. Additionally, it can be seen that the GPU memory is under-utilized for smaller models like LLaMA-2 7B and 13B with ZeRO-3 + LoRA. Consequently, our study shows that for fine-tuning larger models with tens or hundreds of billions of parameters, enabling ZeRO-3 is indispensable. Furthermore, when used in conjunction with LoRA, the overall memory requirements significantly reduce in practice.
@@ 5.4 Long Context Fine-tuning
Figure 3: Impact of varying context length on GPU Memory usage and fine-tuning time with and without Flash-Attention 2 for LLaMA-2 70B. Enabling Flash-Attention 2 on A100s significantly lowers the memory consumption and runtime for larger context lengths such as 4096.
Context length is a critical aspect when fine-tuning LLMs. This is especially true when fine-tuning with datasets comprising of very long sequences of text. Without Flash-Attention 2 (FA2), attention computations scale quadratically in sequence length. Since V100 GPUs do not support FA2, the impact of using longer context lengths can be significant on GPU memory consumption. To study the impact of long context lengths during fine-tuning process, we plot memory and runtime on A100s (with and without FA2 enabled) and V100s (without FA2) across different context lengths. Specifically, we fine-tune LLaMA-2 70B model with context-lengths of 1024, 2048 and 4096 separately on 8xV100 and 8xA100 GPUs. LoRA and Gradient Checkpointing are enabled on top of ZeRO-3 during fine-tuning.
Figures 3 demonstrates the impact of Flash-Attention 2 on GPU memory and runtime. Our results show an expected trend as context length is increased to 4096. There is a significant reduction in GPU memory usage and runtime when using Flash-Attention 2 than without it. As seen in figure 3, this trend is consistent when comparing runs with and without Flash-Attention 2 on two A100 GPUs, or on A100 and V100 GPUs respectively. Our conclusion is as follows. High-performance GPUs like A100s, due to their larger HBMs, can support relatively long context fine-tuning without Flash-Attention 2. However, it is optimal to activate Flash-Attention 2 on compatible GPU architectures as that allows additional memory savings and reduced runtime. This becomes particularly important when training on long text data with models that allow fine-tuning with larger context lengths.
@@ 5.5 Fine-tuning under GPU Resource Constraints
Figure 4: Optimal configurations for fine-tuning LLMs of different sizes when using V100 GPUs. All LLaMA-2 experiments were run using 8xV100s, whereas Falcon-180B required 16xV100s. Flash-Attention 2 is omitted as it is not supported on V100s.
@@
We fine-tune LLaMA-2 7B, 13B, 70B and Falcon-180B model over five different context lengths: [256, 512, 1024, 2048, 4096]. Adhering to the resource constraint specifications, our experiments are done on a Standard_ND40rs_v2 (8xV100 GPUs) with 32 GB of HBM. The only deviation from this setup is for Falcon-180B model. Due to its large size, we use two nodes (16xV100 GPUs) during fine-tuning. Figure 4 illustrates the preferred set of optimizations that should be enabled to meet GPU memory requirements while optimizing for runtime. Note that while there can be other combinations that allow successful fine-tuning, we select those which minimize the fine-tuning time. As an example, users can choose to enable all optimizations, but often they are not all going to be warranted, and may unnecessarily increase the fine-tuning time.
@@
    As context length or model size increases, it becomes important to shift to a higher ZeRO-DP stage to facilitate fine-tuning. This trend is clearly visible in the table, where a move from left to right or top to bottom corresponds to a higher ZeRO-DP stage.
@@
    Gradient Checkpointing (GC) is an effective memory-saving optimization technique, especially for large models. As depicted in figure 4, enabling GC for LLaMA-2 13B and 70B allowed support of context lengths from 512 to 4096, which would have not been possible otherwise. In a similar vein, GC is essential for enabling fine-tuning of Falcon-180B on all our examined context lengths.
Although Flash-Attention 2 was not supported on V100 GPUs, our experimental results suggest that it should always be enabled on supported architectures. It is worth mentioning that the configurations outlined in figure 4 can be applied for fine-tuning on GPUs with higher HBMs than V100s. While these configurations may not be optimal for all GPUs, our findings can be extended to other GPUs with further investigation.

### v2

@@ Abstract
Fine-tuning large language models is a popular choice among users trying to adapt them for specific applications. However, fine-tuning these models is a demanding task because the user has to examine several factors, such as resource budget, runtime, model size and context length among others. A specific challenge is that fine-tuning is memory intensive, imposing constraints on the required hardware memory and context length of training data that can be handled. In this work, we share a detailed study on a variety of fine-tuning optimizations across different fine-tuning scenarios. In particular, we assess Gradient Checkpointing, Low-Rank Adaptation, DeepSpeed’s Zero Redundancy Optimizer and FlashAttention. With a focus on memory and runtime, we examine the impact of different optimization combinations on GPU memory usage and execution runtime during fine-tuning phase. We provide our recommendation on the best default optimization for balancing memory and runtime across diverse model sizes. We share effective strategies for fine-tuning very large models with tens or hundreds of billions of parameters and enabling large context lengths during fine-tuning. Furthermore, we propose the appropriate optimization mixtures for fine-tuning under GPU resource limitations.
@@ 1 Introduction
The remainder of this work is organized as follows. We review the existing fine-tuning optimization literature in section 2. Section 3 presents an overview of commonly used fine-tuning optimizations: Gradient Checkpointing, Low-Rank Adaptation, DeepSpeed’s Zero Redundancy Optimizer and FlashAttention. We then describe in Section 4 the theory behind GPU memory usage and compare theoretical values with our empirical results. In Section 5, we discuss our experimental results and analyze the impact of using fine-tuning optimizations in conjunction with each other. Finally, we share the key conclusions and directions for future work in Section 6.
@@ 2 Related Work
Fine-tuning large language models is a computational and memory intensive process. Several optimizations proposed in existing literature either reduce the memory consumption, or the runtime, or accomplish both during fine-tuning. Tensor and pipeline parallelism are types of model parallelism techniques which enable large models to fit in memory for training. In tensor parallelism (Shoeybi et al., 2020), the model is split tensor-wise among GPUs. Conversely, pipeline parallelism (Huang et al., 2019) distributes the model layer-wise across GPUs. DeepSpeed’s Zero Redundancy Optimizer (ZeRO) (Rajbhandari et al., 2020) integrates data and model parallelism, by partitioning model states across data-parallel processes, which are reconstructed when required during training. Fully Sharded Data Parallel (FSDP) (FairScale authors, 2021) is similar to ZeRO and supports model parameter sharding to train large models.
Contrary to parallelism strategies, Parameter-Efficient Fine-Tuning (PEFT) methods (Ding et al., 2023; Han et al., 2024) modify a small proportion of model parameters, while keeping the others frozen during training. This reduction in total trainable parameters subsequently lowers the memory footprint while still achieving comparable performance to full fine-tuning. Low-Rank Adaptation (Hu et al., 2021) is a widely used PEFT technique, where low rank tensor formulation of model parameters are updated in the training process. Prefix tuning (Li and Liang, 2021) and Prompt tuning (Lester et al., 2021) are examples of PEFT method where trainable embeddings are injected at the model or input level, while the entire model is frozen. PEFT methods can easily integrate with parallelism techniques such as ZeRO, enhancing the gains in terms of memory and runtime reduction during fine-tuning.
Certain optimizations are designed to address specific memory bottlenecks during fine-tuning. Gradient computation during backward pass can consume significant GPU memory for large models because of the need to maintain large number of activations in memory. Gradient Checkpointing (Sohoni et al., 2019) saves only few activations from forward pass, while recomputing most of them during backward pass, to reduce the memory usage at the cost of increased runtime. Attention calculations can form memory bottlenecks when context length is very large. FlashAttention (Dao et al., 2022; Dao, 2023) methods focus on optimizing the memory consumed during attention calculation operations. Both Gradient Checkpointing and FlashAttention can be combined with PEFT and ZeRO.
@@ 3.1 Gradient Checkpointing
3.2 Low-Rank Adaptation
Low-Rank Adaptation (LoRA) reduces the number of trainable parameters which in turn lowers GPU memory requirements during fine-tuning. It achieves this by freezing the pre-trained model weights and subsequently injecting trainable rank decomposition matrices into each of the chosen layers of the Transformer architecture (Hu et al., 2021). The number of trainable parameters are reduced by orders of magnitude, bringing down the cost of fine-tuning, while also preserving the quality of the results. As an example, with LoRA fine-tuning using rank set to 64, the number of trainable parameters for a 70 billion parameter model is reduced to about 131 million parameters ($\sim$0.19of the original model size).
3.3 Zero Redundancy Optimizer
DeepSpeed’s Zero Redundancy Optimizer (ZeRO) (Rajbhandari et al., 2020) is a memory optimization technique that provides the benefits of model and data parallelism, while alleviating the limitations of both. ZeRO-powered data parallelism (ZeRO-DP) partitions the model states – parameters, gradients, and optimizer states – across the data-parallel processes and uses a dynamic communication schedule to share the necessary model states across processes. ZeRO-DP provides three optimization stages, which successively provide larger and larger memory reduction while incurring some runtime overhead. ZeRO-DP based fine-tuning can achieve memory reduction from 4 to 8 times for Stages 1 and 2 respectively, and all the way up to linear memory reduction for Stage 3 (given a large number of GPUs). This does come at the expense of runtime, especially for ZeRO-DP Stage 3. ZeRO optimization is further enhanced by the inclusion of ZeRO-Offload and ZeRO-Infinity (Ren et al., 2021; Rajbhandari et al., 2021), which enable offloading the optimizer states and model parameters to CPU respectively. While ZeRO-Offload is available across all three ZeRO-DP stages, ZeRO-Infinity is available only for stage 3. Without using DeepSpeed’s ZeRO kind of optimizations, many large models (several billions of parameters) are practically impossible to train using just a handful of V100 or A100 GPUs.
@@ 3.4 FlashAttention
FlashAttention helps achieve attention calculations in linear instead of quadratic complexity with respect to the sequence/context length (Dao et al., 2022). It leverages tiling and recomputation techniques to significantly speed up attention computation. It makes judicious use of the Static Random Access Memory (SRAM), which is the most expensive but smallest sized memory unit within the GPU, minimizing reads and writes between SRAM and HBM. In FlashAttention-2 (FA2), there are further optimizations to reduce the slower non-matrix multiplication operations and to parallelize forward and backward passes along the sequence length dimension (in addition to the batch and number of heads dimensions) (Dao, 2023).
@@ 4 Theoretical Analysis of GPU Memory Requirements
Table 1: Comparison between theoretical and empirical GPU memory allocated (in GB) across the three stages of ZeRO-DP. Fine-tuned model is Llama 2 7B using 8xA100 GPUs (each with 80 GB of HBM). The theoretical estimates can be used to approximate the empirical memory requirements before running the fine-tuning jobs.
Table 1 shows our results of comparing the theoretically expected and empirical GPU memory allocated for fine-tuning Llama 2 7B across ZeRO-DP stages. We use the notation ZeRO-1, ZeRO-2 and ZeRO-3 to denote ZeRO-DP stages 1, 2 and 3 respectively. The theoretical memory is computed by summing the memory from model states and activations. We use the following model default parameters: number of transformer layers=32, hidden dimension size $h$=4096 and attention heads $a$=32 to theoretically compute the total activation memory of 1.48 GB. The empirical GPU memory refers to the peak GPU memory allocated during fine-tuning. We show empirical memory with and without CPU offloading of optimizer states. The experimental setup details are described in subsection 5.1.
@@ 5 Experiments
We begin this section by analyzing the impact of fine-tuning optimizations on memory and runtime. The results of our analysis are used to recommend a set of optimizations that can act as a balanced default to optimize for memory and runtime during fine-tuning. Subsequently, we delve into fine-tuning large models (with tens of billions of parameters) and explore the optimizations which make such fine-tuning possible. Following this, we examine the role of FlashAttention-2 during fine-tuning of LLMs on long context data. Lastly, we investigate the specific case of fine-tuning with resource-constrained GPUs, specifically V100s. Our goal is to identify the fine-tuning optimizations that can enable efficient fine-tuning across various model sizes and context lengths, even with limited resources.
@@ 5.1 Setup
We fine-tune models from Llama 2 (7B, 13B, 70B) (Touvron et al., 2023) and Falcon (180B) (Almazrouei et al., 2023) families on Causal Language Modeling task. Following are the details of our experimental setup:
@@ 5.2 Best Default Optimizations to Balance Memory and Runtime
Figure 2: GPU memory usage and fine-tuning runtime for different optimization configurations across ZeRO-1, ZeRO-2 and ZeRO-3 for Llama 2 7B. Using LoRA with ZeRO-2 provides the best balance between memory usage and runtime.
We have already established that for a relatively small 7B parameter model, it is infeasible to fine-tune without any optimizations, even when using A100s. Therefore, we use DeepSpeed’s ZeRO which provides both model and data parallelism. We examine the set of fine-tuning optimizations from section 3 that should be enabled by default in a fine-tuning framework. Our criteria for selecting such optimizations is based on their ability to strike an optimal balance between memory usage and runtime in conjunction with ZeRO-DP.
We fine-tune Llama 2 7B on one node (8xA100 GPUs) with five distinct configurations: (a) no optimizations among GC, LoRA and FA2 enabled (b) with GC (c) with LoRA (d) with FA2 and (e) without CPU offloading. Each configuration is enabled with all three ZeRO-DP stages. We chose A100 GPUs for these experiments due to their substantial memory capacity of 80 GB. This ample memory allows us the flexibility to experiment with various optimizations by toggling them on and off, while avoiding out-of-memory failures.
@@ 5.3 Fine-tuning Large Models
Fine-tuning large models such as Llama 2 70B or Falcon 180B can often run into out-of-memory (OOM) errors if the right set of optimizations are not enabled. Enabling ZeRO-3 is a viable alternative in such scenarios though it comes at the expense of increased runtime. Let’s examine the model state memory term in ZeRO-3, fine-tuned over $M$ nodes with $N$ GPUs per node:
@@
Llama 2 Model GPU Memory Usage (GB) Time (seconds)
@@
Table 2: GPU memory usage and fine-tuning time for Llama 2 Models with ZeRO-3 + LoRA on 8xV100 GPUs.
We were able to fine-tune Llama 2 70B using a combination of ZeRO-3 + LoRA on a single node with 8xV100 GPUs (32 GB HBM). Table 2 shows the actual GPU memory consumed to be around 15.54 GB for the 70B model in our experiments. Additionally, it can be seen that the GPU memory is under-utilized for smaller models like Llama 2 7B and 13B with ZeRO-3 + LoRA. Consequently, our study shows that for fine-tuning larger models with tens or hundreds of billions of parameters, enabling ZeRO-3 is indispensable. Furthermore, when used in conjunction with LoRA, the overall memory requirements significantly reduce in practice.
@@ 5.4 Long Context Fine-tuning
Figure 3: Impact of varying context length on GPU memory usage and fine-tuning time with and without FlashAttention-2 for Llama 2 70B. Enabling FlashAttention-2 on A100s significantly lowers the memory consumption and runtime for larger context lengths such as 4096.
Context length is a critical aspect when fine-tuning LLMs. This is especially true when fine-tuning with datasets comprising of very long sequences of text. Without FlashAttention-2 (FA2), attention computations scale quadratically in sequence length. Since V100 GPUs do not support FA2, the impact of using longer context lengths can be significant on GPU memory consumption. To study the impact of long context lengths during fine-tuning process, we plot memory and runtime on A100s (with and without FA2 enabled) and V100s (without FA2) across different context lengths. Specifically, we fine-tune Llama 2 70B model with context-lengths of 1024, 2048 and 4096 separately on 8xV100 and 8xA100 GPUs. LoRA and Gradient Checkpointing are enabled on top of ZeRO-3 during fine-tuning.
Figure 3 demonstrates the impact of enabling FA2 on GPU memory and runtime. Our results show an expected trend as context length is increased to 4096. There is a significant reduction in GPU memory usage and runtime when using FA2 than without it. As seen in Figure 3, this trend is consistent when comparing runs with and without FA2 on two A100 GPUs, or on A100 and V100 GPUs respectively. Our conclusion is as follows. High-performance GPUs like A100s, due to their larger HBMs, can support relatively long context fine-tuning without FlashAttention-2. However, it is optimal to activate FlashAttention-2 on compatible GPU architectures as that allows additional memory savings and reduced runtime. This becomes particularly important when training on long text data with models that allow fine-tuning with larger context lengths.
@@ 5.5 Fine-tuning under GPU Resource Constraints
Figure 4: Optimal configurations for fine-tuning LLMs of different sizes when using V100 GPUs. All Llama 2 experiments were run using 8xV100s, whereas Falcon 180B required 16xV100s. FlashAttention-2 is omitted as it is not supported on V100s.
@@
We fine-tune Llama 2 (7B, 13B, 70B) and Falcon (180B) models over five different context lengths: [256, 512, 1024, 2048, 4096]. Adhering to the resource constraint specifications, our experiments are done on a Standard_ND40rs_v2 (8xV100 GPUs) with 32 GB of HBM. The only deviation from this setup is for Falcon 180B model. Due to its large size, we use two nodes (16xV100 GPUs) during fine-tuning. Figure 4 illustrates the preferred set of optimizations that should be enabled to meet GPU memory requirements while optimizing for runtime. Note that while there can be other combinations that allow successful fine-tuning, we select those which minimize the fine-tuning time. As an example, users can choose to enable all optimizations, but often they are not all going to be warranted, and may unnecessarily increase the fine-tuning time.
@@
    As context length or model size increases, it becomes important to shift to a higher ZeRO-DP stage to facilitate fine-tuning. This trend is clearly visible in Figure 4, where a move from left to right or top to bottom corresponds to a higher ZeRO-DP stage.
@@
    Gradient Checkpointing (GC) is an effective memory-saving optimization technique, especially for large models. As depicted in Figure 4, enabling GC for Llama 2 13B and 70B allowed support of context lengths from 512 to 4096, which would have not been possible otherwise. In a similar vein, GC is essential for enabling fine-tuning of Falcon 180B on all our examined context lengths.
Although FlashAttention-2 was not supported on V100 GPUs, our experimental results suggest that it should always be enabled on supported architectures. It is worth mentioning that the configurations outlined in Figure 4 can be applied for fine-tuning on GPUs with higher HBMs than V100s. While these configurations may not be optimal for all GPUs, our findings can be extended to other GPUs with further investigation.